OpenStudy (anonymous):

a rectangle has a perimeter 20m. express the area of the rectangle as a function of the length of one of its sides.

OpenStudy (anonymous):

Let L, H be the length and the height of the rectangle. So the perimeter is 2L+2H = 20 i.e. L+H=10 => H = 10 - L Area of the rectangle given by L X H You can substitute H in above equation to find the area in terms of L